OpenStudy (anonymous):

How do you solve radical 5 over radical 3 ...and 3 radical 12 times radical 6

OpenStudy (anonymous):

Define 'solve'.

OpenStudy (anonymous):

1. Multiply top and bottom by sqrt 3. The numerator will be sqrt 5 * sqrt 3 which can be combined to sqrt (5*3) = sqrt (15). The denominator will be just 3. So the answer is sqrt(15) / 3. 2. You can rewrite the sqrt 12 / sqrt 6 as sqrt(12/6) which = sqrt 2.. So the answer is 3*sqrt(2).

OpenStudy (anonymous):

lmao i got it right the first time and i change it and got it wrong, omg drinking too much

OpenStudy (anonymous):

oh great and i hit delete

OpenStudy (anonymous):

don't drink and derive...